The length of a rectangular solid is twice the width, and the height is three times the width. If the volume is 162 cubic meters, find the surface area of the solid.
Answer by ramkikk66(644)

Let the width be x
Then length = 2x
Height = 3x
Volume = Length * width * height = x*2x*3x = 6*x^3 = 162
x^3 = 162/6 = 27
x = 3
Width = 3
Length = 6
Height = 9
:)